15liters of 96% alcohol added 20 liters of water,what is the alcohol %?
Answer by ewatrrr(24785)

 
Hi
water = water
.04(15L) + 20L = P(15L+20L)
20.6/35 = P (of water) = .5886
alcohol: 1 - .5586 = .4114 0r 41.44%